OpenStudy (lovelyanna):

Simplify. (x^2 + 5x + 8)(x^2 + 2x + 6) A. 2x^2 + 7x + 14 B. x^4 + 7x^3 + 24x^2 + 46x + 48 C. x^4 + 2x^3 + 6x^2 + 7x + 14 D. x^4 + 10x^2 + 48

OpenStudy (ranga):

(x^2 + 5x + 8)(x^2 + 2x + 6) = x^2(x^2 + 2x + 6) + 5x(x^2 + 2x + 6) + 8(x^2 + 2x + 6) = ? multiply each term within the parenthesis with the term on the outside. Gather like terms and simplify.

OpenStudy (lovelyanna):

Cross out A, correct?

OpenStudy (ranga):

correct.

OpenStudy (lovelyanna):

I don't think C is the answer either..?

OpenStudy (ranga):

Don't guess. The multiplication is not hard to do.

OpenStudy (ranga):

x^2(x^2 + 2x + 6) = x^4 + 2x^3 + 6x^2 5x(x^2 + 2x + 6) = ? 8(x^2 + 2x + 6) = ? Add up the right hand side and simplify.

OpenStudy (lovelyanna):

B?

OpenStudy (lovelyanna):

This is confusing :/

OpenStudy (ranga):

x^2(x^2 + 2x + 6) = x^4 + 2x^3 + 6x^2 5x(x^2 + 2x + 6) = + 5x^3 + 10x^2 + 30x 8(x^2 + 2x + 6) = + 8x^2 + 16x + 48 -------------------------------------- Add: x^4 + 7x^3 + 24x^2 + 46x + 48

OpenStudy (lovelyanna):

Uhhh... I don't get this :/ Easier explanations please?

OpenStudy (ranga):

(x^2 + 5x + 8) * (x^2 + 2x + 6) Take the first term x^2. Multiply every term within the second parenthesis by x^2. You get: x^4 + 2x^3 + 6x^2. Is this part clear or do you have questions?

OpenStudy (lovelyanna):

So... C is the answer?

OpenStudy (ranga):

The answer was given in my reply before last: x^4 + 7x^3 + 24x^2 + 46x + 48

OpenStudy (lovelyanna):

Oh, I did say B. Thanks!

OpenStudy (ranga):

I did not say it was wrong. But rather than just calling out each answer choice it is better to do the math as the multiplication here is not difficult. You are welcome.